It is approximately 300 miles from Chicago, Illinois, to St. Louis, Missouri. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
a) How far have you traveled after 3 hours?
Answer: 180 miles
Show work in this space: 3(60) = X (I think this is how it has to be done)

b) How far have you traveled after 4 hours?
Answer:4(60)
Show work in this space:
c) How far have you traveled after t hours? i.e. write a linear function that expresses the distance traveled, d, as a function of time, t.
Answer:

d) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled 3 hours?
Answer:
Show work in this space:
e) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled 4 hours?
Answer:
Show work in this space:

f) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled t hours? i.e. write a linear function that expresses the distance to be traveled to reach St. Louis, s, as a function of time, t.
Answer:2 more hours

Answer by 303795(595)

It is approximately 300 miles from Chicago, Illinois, to St. Louis, Missouri. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
a) How far have you traveled after 3 hours?
Answer: 180 miles
Show work in this space: 3(60) = X (I think this is how it has to be done)
b) How far have you traveled after 4 hours?
Answer: 240 miles
Show work in this space: 4 times 60 = 240 miles
c) How far have you traveled after t hours? i.e. write a linear function that expresses the distance traveled, d, as a function of time, t.
Answer: d = 60t
The distance is found by multiplying the time in hours by the speed so after t hours you have travelled distance = 60t
d) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled 3 hours?
Answer: 120 miles
Show work in this space: After travelling for 3 hours you have travelled 180 miles (see question a) so you have 300 - 180 = 120 miles to go
e) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled 4 hours?
Answer: 60 miles
Show work in this space: After travelling for 4 hours you have travelled 240 miles (see question b) so you have 300 - 240 = 60 miles to go
f) How far will you HAVE LEFT to travel to reach St. Louis after you have traveled t hours? i.e. write a linear function that expresses the distance to be traveled to reach St. Louis, s, as a function of time, t.
Answer: s = 300 - 60t
Show work in this space: After travelling for t hours you have travelled 60t miles (see question c) so you have 300 - 60t miles to go